1. Two lighthouses A and B are 25km apart and A is due west of B. A submarine S is on a bearing of 137degrees from A and on a bearing of 170degrees from B. Find the distance of S from A and the distance of S from B.


2. In the triangle PQR, PQ= 7cm, PR= 8cm and QR= 9cm. Find angle QPR.\


3. Find the smallest angle in a triangle whose sides are of length 3x, 4x and 6x.


4. Shop Y is 9km due North of shop Z. Shop X is 8km from Y, 5km from Z and somewhere to the west of line YZ.

a) Draw the triangle XYZ and find angle YZX.

b) To improve sales, shop X is moved due to South until it is due West of Z. Find how far it moved.


1
Expert's answer
2021-08-11T13:29:22-0400

1.


Using Sine rule:


"\frac{a}{sin\left(A\right)}=\frac{b}{sin\left(B\right)}"


"\frac{25}{sin\:33}=\frac{b}{sin\:47}"


"b=\frac{25\times sin\:47}{sin\:33}"


"b=33.57km"


Hence, distance of S to B is 33.57km.


"\frac{b}{sin\left(B\right)}=\frac{c}{sin\left(C\right)}"


"\frac{33.57}{sin\left(47\right)}=\frac{c}{sin\left(100\right)}"


"c=\frac{33.57\times sin\:100}{sin\:47}"


"c=45.20km"


Hence, the distance of S to A is 45.20km.


2.


To find angle QPR we use the cosine rule:

"a^2\:=b^2+c^2-2bcCosA"


"Cos\:A=\frac{b^2+c^2-a^2}{2bc}"


"Cos\:A=\frac{7^2+8^2-9^2}{2\times 7\times 8}"


"Cos\:A=0.2857"


"A=Cos^{-1}\left(0.2857\right)"

"A\approx 73.4^o"


Hence angle QPR is "73.4^o"


3.

"Angle\:1=3x"

"Angle\:2=4x"

"Angle\:3=6x"


"Angle\:sum\:property\:of\:a\:triangle=180"

"\:\:3x+6x+4x=180"

"13x=180"

"x=13.85"


The smallest angle will be: "3x"

"3x=3\times13.85\approx41.5^o"


4.


(a). The angle YZX can be determined by cosine rule:

"a^2\:=b^2+c^2-2bcCosA"


"Cos\:A=\frac{b^2+c^2-a^2}{2bc}"


"Cos\:A=\frac{5^2+9^2-8^2}{2\times 5\times 9}"


"A=Cos^{-1}\left(0.4667\right)"


"A\approx62.2^o"


Hence, angle YZX is "62.2^o"


(b).


To determine the distance Shop X moved to a reach a new location we use the formula:

"Distance=\frac{\theta }{360}\:\times 2\times \pi \times r\:"


"Distance=\frac{207.8^o}{360^o}\:\times \:2\times \:\frac{22}{7}\:\times \:5"


"Distance\approx 18.14km"


Hence, shop X moved 18.14 km to reach a new location.